OpenStudy (anonymous):

Solve: -1/2x + 1 = −x + 8

OpenStudy (lgbasallote):

\[\Large -\frac 12x + 1 = -x + 8\] \[\Large -\frac 12x + 1 = 8-1\] \[\Large \frac 12 x = 7\] \[\Large x = 7 \times 2\] \[\Large x = 14\] does that help?